### v1

# Circumgalactic Enrichment by a Multi-Galaxy Merger at Redshift 6.7

Abstract
Recent JWST observations have uncovered high-redshift galaxies characterized by multiple star-forming clumps, many of which appear to be undergoing mergers [1, 2, 3, 4]. Such mergers are powerful drivers of galaxy evolution, triggering intense star formation [5, 6, 7] that accelerates the formation of massive quiescent galaxies [8]. Mergers can also expel and redistribute the metal-enriched gas from galaxies into their circum-galactic medium (CGM), regulating the chemical component of their environments [9, 10, 11] and creating channels of ionized gas through which ionizing radiation escapes [12, 13]. Here, we report a merger of at least five galaxies, dubbed JWST’s Quintet (JQ), at redshift 6.7, discovered in the JWST GOODS-South field. This system resides in a small area ( pkpc2), containing over 17 galaxy-size clumps with a total stellar mass of . The JQ system has a total star formation rate of 240 – 270 yr-1, placing it dex above the median star formation rate–mass main sequence at this epoch. The high mass and star formation rate of the JQ galaxies are sufficient to produce massive, inactive galaxies identified at later times, by redshift 4 [14, 15, 16], suggesting this may be a progenitor event for such systems. We detect a large [O III]+H emitting gaseous halo surrounding and connecting four galaxies in the JQ, consistent with the interaction between them. This indicates that the CGM of JQ is already enriched by the heavy elements generated from the galaxies through merger-induced tidal stripping or galactic outflows. Our findings offer a plausible evolutionary pathway for the formation of massive quiescent galaxies observed at lower redshifts and provide direct evidence for the metal enrichment of their environments through galaxy mergers, just 800 Myr after the Big Bang.
We identified a group of five emission line galaxies (ELGs) at redshift 6.7, which we call JWST’s Quintet (JQ), from a study of ELGs based on imaging and spectroscopic data from the JWST Advanced Deep Extragalactic Survey [17] (JADES). JADES, one of the deepest JWST/NIRCam imaging surveys to date, provides continuous wavelength coverage from 0.8 to 5.0 m across 14 broad- and medium-band filters (see Methods). The JADES multi-wavelength imaging data enables precise photometric redshift estimation with an uncertainty of [17], allowing reliable distances to galaxies and distinguishing galaxies in close physical proximity. We identified ELGs with strong [O III]+H nebular emission as well-detected () sources with photometric redshifts in the range 6.7 to 7.6 with a excess of flux emission in the F410M medium-band image compared to the F444W broad-band image of JADES. At these redshifts, the [O III] and H emission lines fall in the F410M bandpass, and the flux excess criteria identities objects with EW([O III]+H) 300 Å. The spatial distribution of ELGs revealed an overdensity of the five JQ galaxies within a region, see Figure 1, corresponding to proper kpc2 (pkpc2) Of these, two galaxies (ELG1 and ELG5) have spectroscopically confirmed redshifts of 6.707 and 6.712 based on the NIRSpec spectroscopy [18]. These two galaxies reside at the head and the tail of the JQ with a projected distance of 18.6 pkpc and a velocity difference of 195 km s-1. The remaining three ELGs (ELG2, ELG3, ELG4) reside between them and have similar spectral energy distributions (SEDs) as ELG1 and ELG5 with consistent photometric redshifts (see Figure 2). This is evidence all five ELGs are in the same physical structure.
The JQ appears to be a high-redshift analog of a well-known galaxy group, Stephan’s Quintet, at a distance of 85 Mpc. Though, unlike Stephan’s Quintet, where only four of the five galaxies are physically associated and interacting in mergers [19, 20], we believe all galaxies in JQ are physically associated. Furthermore, there is evidence that the entire structure of ELG1, ELG2, ELG3, and ELG4 is connected by a diffuse emission of [O III]+H only seen in the F356W image. At the redshift of JQ, the [O III] line falls within the F410M band, facilitating the detection of this system, but the F356W band encompasses both [O III] and H with much higher transmission than F410M. We therefore use F356W for analysis in this paper. The implication is that these galaxies have undergone an interaction that produced the diffuse [O III]+H emission. In ELG2, a notably extended tail (labeled C5 in Figure 1) points toward ELG1 in the projected plane, with several faint star-forming clumps (C4, C12, C13; clump identification is described below) distributed along it. These clumps likely trace the star formation induced by the galaxy mergers. ELG3 also shows a disturbed morphology with two clumps (C8-1, C8-2) connected through diffuse emission to ELG4. While we do not detect diffuse [O III]+H emission nor star-forming clumps between ELG5 and ELG3, the small velocity offset of 195 km s-1 between ELG5 and ELG1 and the small projected distance of 6.2 pkpc between ELG5 and ELG3 strongly suggests that ELG5 is also interacting with the other ELGs.
The JQ galaxies show a high frequency of spatially resolved clumps. We adopt an automatic clump identification algorithm combined with visual inspection to detect the clumps in this structure. We measure the ELG sizes and photometry by fitting their surface brightness with Sérsic profile models, and we measure the galaxy properties by fitting models to the SEDs. We identified 17 clumps associated with the JQ and 7 foreground galaxies. Extended Data Table 1 lists the properties of the JQ clumps. To account for the diffuse emission surrounding the clumps, we also perform a spatially resolved SED fitting to measure the integrated properties of the entire system. We derive a total stellar mass of and star formation rate (SFR) of yr-1. We remove the foreground galaxies from the images based on their best-fit Sérsic profiles and present a clean map of JQ in Figure 1c,d. See Methods for details.
The sizes and SFRs of the JQ clumps are significantly larger than those of the star clusters and star cluster-size clumps within individual galaxies resolved through the lensing technique [21, 22, 23]. Their sizes, stellar masses, and SFRs are more consistent with the high-redshift spatially unresolved field galaxies observed with JWST NIRCam [24]. Although recent JWST observations have reported numerous galaxy systems containing such galaxy-size clumps [4, 2, 3, 1], JQ stands out as the first system identified with over 10 galaxy-size clumps. This high density of galaxy-size clumps within a single galaxy system is extremely rare and challenges predictions from cosmological simulations. For example, work focusing on clump formation during mergers between high-redshift galaxies does not identify any systems with more than four clumps at [25]. Comparing JQ to predictions from the 6.5 – 7.0 galaxy systems in a 2-deg2 simulation [26], we identify only 8 systems with galaxy-size clumps out from a parent sample of 29,696 systems. Furthermore, from this simulation, we would expect only 0.26 JQ-like systems in the JADES survey at – 7.6. This indicates the number density of galaxy-size clumps in the JQ is higher than galaxy systems in cosmological simulation, and is a rare configuration. The small area of JADES F410M image ( arcmin2) implies an observed comoving number density of cMpc3 for JQ, almost quadruple the expected number density from the cosmological simulation ( cMpc3). see Methods for details.
The clumps associated with the JQ galaxies show evidence of elevated SFRs compared to other galaxies at these epochs. Figure 3 compares the SFRs of the clumps to the galaxies at similar redshift and the SFR–Mass Main Sequence (SFMS) using results from the literature [27, 29, 30, 28]. In all cases, the clumps in JQ with stellar masses have SFRs more than 1 dex above the measured SFMS relation, indicating they are in a bursting phase [27]. The best-fit model star-formation histories of these clumps show that they could form all of their stellar mass in the previous 10 Myr, but this is a lower limit as the recently formed stars could outshine previous generations of stars [32, 33]. Nevertheless, the short formation timescales suggest that the star formation in these clumps was recently triggered by major mergers in JQ [25]. The clumps with higher stellar mass, (C3, C8, C9) have SFRs consistent with the median SFR–Mass relation at this epoch. The total SFR of galaxies in the JQ system is higher than the SFMS by dex. This indicates that the elevated SFR and the formation of clumps is correlated with the major merger in the JQ system.
The elevated SFR triggered by major mergers in JQ provides a plausible evolutionary pathway for forming massive quiescent galaxies, recently discovered at 4.5 – 5. The existence of these objects has challenged our understanding of rapid galaxy growth and quenching over short timescales [14, 16, 31]. The high stellar mass and SFR of the JQ system are consistent with those needed to produce massive (), quiescent galaxies observed at redshift 4 – 5 [14, 16] (see Figure 4). The expected merging timescale for the JQ system is – 350 Myr (see Methods). If the JQ galaxies sustain the current SFR during this period, they will build a single massive galaxy with a stellar mass of SFR – . Simulations show that such a major merger event is expected to rapidly consume cold gas and trigger strong feedback from black hole activities, ultimately quenching star formation in the JQ galaxies [34, 35, 36]. Furthermore, the observed number density of JQ-like systems is consistent with that of massive quiescent galaxies at redshift , estimated to be Mpc-3 [16, 14]. These factors suggest that JQ is a likely candidate for a progenitor of the massive quiescent galaxies identified at redshifts 4 – 5. To test this interpretation requires a detailed study of the kinematics of the individual JQ clumps, which is possible with future JWST observations.
The NIRCam F356W image of JQ is dominated by the redshifted [O III]+H emission line. The image shows extended, diffuse emission, compared to the NIRCam F335M and F277W images, which is evidence of a [O III]+H halo surrounding the JQ system. We estimate the amount of [O III]+H emission by subtracting the continuum contribution from the F356W image (see Methods) and present the [O III]+H halo as the orange contours in Figure 1 (the original [O III]+H map is shown in Extended Data Figure 3). The [O III]+H halo surrounds the entirety of the ELG1, ELG2, ELG3, and ELG4 galaxies, extending more than 20 kpc. The total [O III]+H luminosity is erg s-1. Notably, we observe a “bridge” in the [O III]+H emission connecting galaxies ELG1 and ELG2. This bridge extends to 5 kpc and includes the clumps C12 and C13. The presence of this large-scale [O III]+H halo requires oxygen-enriched gas in the circum-galactic medium (CGM), ejected and/or stripped efficiently from the galaxies in JQ, particularly from ELG1 and ELG2, as evidenced by the bright [O III]+H bridge connecting them.
Ejection of gas is observed as galactic outflows driven by massive stars, supernovae, or active galactic nuclei (AGN). This has been proposed to explain the origin of [O III]+H halos and CGM enrichment in galaxies in the distant universe [37, 38, 3]. The large [O III]+H halos of the size seen in JQ have only been observed in the quasars and AGNs [38, 39, 40]. There is currently no evidence of AGN in the JQ galaxies, particularly in ELG1 and ELG2 given the spatially resolved morphology of their clumps. There is also no evidence of any broad emission component in the existing NIRSpec M-grating spectrum of ELG1, indicating if such an outflow exists, it is not powerful enough to eject large amounts of oxygen-enriched gas to CGM, at least along the line of sight. We compared the properties of outflows in a state-of-the-art model and found the outflows produced by supernovae in ELG1 and ELG2 and their adjacent clumps are not sufficient to explain the observed luminosity of the [O III]+H “bridge” between the two galaxies (see Methods). Consequently, the outflows are unlikely to be the dominant driver of CGM enrichment in JQ.
It is therefore more plausible that the [O III]+H halo of JQ results from oxygen-enriched gas stripped out of the galaxies through interactions and tidal forces [11, 41]. The stripped gases are shock-heated in galaxy collisions [7, 42], and dispersed to large radii leading to large hot gaseous halos [43]. This scenario is believed to be responsible for the diffuse emission halo in Stephan’s Quintet, where the shocked-heated gas has a temperature of K and appears as a ridge of X-ray and radio synchrotron emission [44, 45], residing between the two colliding galaxies. Such emission is below the sensitivity of X-ray and radio telescopes given JQ’s distance, but it is possible to confirm the shock-heated origin of the gas using excitation diagnostics of emission lines, as those have been performed in Stephan’s Quintet [46, 47, 48]. Similar to the distribution of shock-heated gas in Stephan’s Quintet, the majority of diffuse [O III]+H emission in JQ also arises from the “bridge” between two merging galaxies, ELG1 and ELG2. Therefore, the evidence supports the interpretation that the [O III]+H halo in JQ results from shock-heated gas, stripped from the galaxies through interactions from a merger.
The discovery of the JQ system at in the JADES data suggests that major mergers of multiple galaxies might be more frequent in the early universe than expected from theory. These mergers can efficiently eject metal-enriched gas into the CGM, thereby contributing to the development of the galactic ecosystem. The elevated star formation associated with mergers offers a unique solution to the tension between the observation and cosmological simulation on the observed overabundance of the massive galaxies and massive quiescent galaxies recently unveiled by JWST.
Methods
Throughout this study, we adopt the Planck 2018 cosmological parameters [49]: , and km s-1 Mpc-1, where and are the densities of total matter and dark energy and is the Hubble constant.
HST and JWST Imaging data
We utilize the published JWST images and spectroscopy from the JWST Advanced Deep Extragalactic Survey (JADES) DR3 [50, 51, 18] to measure the flux densities and to detect clumps in the JQ galaxies. JADES is one of the deepest NIRCam surveys, providing 14 broad- or medium-bands (broad- and medium-band imaging from F090W, F115W, F150W, F182M, F200W, F210M, F277W, F335M, F356W, F410M, and F444W filters from JADES [2], combined with medium-band imaging in F430M, F460M, and F480M [52]) covering wavelengths from 0.8 to 5.0 m. The JADES photometric catalog provides matched-aperture flux densities for the galaxies, and photometric redshifts measured from the SEDs [17, 51, 2].
In addition to the JWST images, we also utilize the archival Hubble Space Telescope (HST) images to complement the JWST images in the UV-to-optical wavelength range and to remove the foreground galaxies. The archival HST images were obtained from several HST observation programs, including CANDELS, HUDF, UVCANDELS, etc [53, 54, 55, 56]. We adopt the reduced HST images provided by the JWST FRESCO collaboration [57], which has already been registered to match the JWST pixel scale (0.03 arcsec).
JWST Spectroscopy data
The spectroscopic data of ELG1 and ELG5 used in this work were also obtained as part of the JADES NIRSpec observations. The JADES program opened three adjacent shutters for each target and nodded between these shutters (perpendicular to the dispersion direction) to improve the background subtraction and bad pixel rejection. JADES obtained low-resolution spectra () using the NIRSpec disperser/filter configuration of PRISM/CLEAR, which provides wavelength coverage from 0.6 to 5.3 m. The exposure times of ELG1 and ELG5 were 2.1 and 9.3 hours, respectively. JADES also obtained medium-resolution spectra () using the disperser/filter configurations of G140M/F100LP, G140M/F070LP, and G235M/F170LP, G395M/F290LP, which provides a full wavelength coverage from 0.7 to 5.3 m. The exposure times of ELG1 for the three configurations were 3.4, 2.8, and 2.8 hours, and the exposure times of ELG5 for the three configurations were 2.3 hours. The JADES NIRSpec data used in this work was downloaded from the MAST HLSP archive111https://archive.stsci.edu/hlsp/jades. Please refer to ref [58] for details.
Photometric redshifts of ELGs
Based on the JADES photometry, we use the Python package Bagpipes [59] to perform SED fitting including measuring redshifts of ELG2, ELG3, and ELG4. We employ Binary Population and Spectral Synthesis v2.2.1 [60, 61] (BPASS) with a broken power-law initial mass function with slopes of for stars with 0.1 – 0.5 and for 0.5 – 100 (model ‘135_100’). For the star formation history parameterization, we use the Gaussian Process model from DENSEBASIS [62], where the star formation history is split into 4 dynamically adjusted time bins, and during each time bin, of the total stellar mass is formed. The metallicity is allowed to span 0.001 – 1 . We adopt the Calzetti [63] dust attenuation law with ranging from 0.0 to 4.0 mag. We include nebular emission with the metallicity of the gas equal to that of the stellar populations, and an ionization parameter, in the range to . We use the uniform priors for the Bagpipes parameters across the allowed range. For the redshift, we also adopt a uniform prior across the range of 5 – 8.5. Figure 2 shows the probability distribution function of the redshifts and the best-fit models for ELG2, ELG3, and ELG4.
We also use Bagpipes to estimate the slitlosses of NIRSpec spectroscopy of ELG1 and ELG5. We adopt a multiplicative scaling factor (assumed to be a second-order Chebyshev polynomial) in Bagpipes to correct the wavelength-dependent spectral flux calibration to the broadband photometry. We allow the zeroth order to vary from 0.1 – 10 because of the relatively large slitloss in ELG1, while the first and second order are allowed to vary from -0.5 – 0.5. In Figure 2, we present the flux-calibrated spectra of ELG1 and ELG5.
| Galaxy | clump | SFR10Myr | MUV | ||||
| ID | ID | yr-1 | [erg s-1] | pkpc | |||
| ELG1 | C1 | ||||||
| C2 | |||||||
| C3 | |||||||
| C4 | |||||||
| ELG2 | C5 | ||||||
| C6 | |||||||
| C7 | |||||||
| ELG3 | C8-1 | ||||||
| C8-2 | |||||||
| C9 | |||||||
| ELG4 | C10 | ||||||
| ELG5 | C11 | ||||||
| C12 | |||||||
| C13 | |||||||
| C14 | |||||||
| C15 | |||||||
| C16 | – | ||||||
| JQ |
Clump identification and photometry
Our clump identification method is illustrated in the Extended Data Figure 1. We identify the clumps from the NIRCam F115W image, as it offers the highest spatial resolution compared to the NIRCam images at longer wavelengths and is unaffected by the Ly break, unlike F090W. We first extract a cutout, corresponding to a physical size of pkpc2. To suppress the outshining components and enhance the visibility of clump cores, we smooth the cutout using a two-dimensional Gaussian kernel ( 1 pixel; 0.03 arcsec) and subtract the smoothed image from the original F115W cutout to produce a contrast image. We then use the Python Package sep [64, 65] to detect clumps in the contrast image and extract their positions. The red circles in Extended Data Figure 1c,d mark the identified clumps. Additionally, we notice several clumps that are clearly detected in the F356W image, but are only marginally detected in F115W. These clumps are manually added to the catalog based on their positions in the F356W image, and they are marked as green circles in Extended Data Figure 1d.
We measure the clump photometry by modeling each clump with a Sérsic profile convolved with the corresponding point spread function (PSF) for each band using the Python package Pysersic [66]. Pysersic requires a set of priors, including clump positions, fluxes, and Sérsic parameters to initialize the fitting. It applies the No-U Turn (NUTS) sampler [67] to simultaneously fit all clumps and produce the posterior distributions. The PSF used in Pysersic is measured using the compact, isolated stars identified in the JADES NIRCam images in each filter.
We first perform the morphological fitting to the F115W image. We adopt the uniform priors for the Sérsic parameters and the Gaussian priors for the positions with the mean to be those measured by sep. Extended Data Figure 1 panels and show the best-fit model and the residual map of F115W, respectively. After obtaining the best-fit parameters from F115W, we use these measurements as the priors for fitting the clumps in other bands. Due to the substantial PSF variation across JWST images (from in F115W to in F444W), deblending adjacent clumps becomes increasingly challenging at longer wavelengths. Therefore, to address this, we fix the pixel centers and Sérsic parameters to the values obtained from F115W, and only the fluxes are allowed to vary when fitting the other bands (F090W, F150W, F200W, F277W, F335M, F356W, F410M, and F444W).
Photometric redshifts and physical properties of clumps
Using the photometry described in the previous section, we perform model fitting for the clumps again using Bagpipes. This step is necessary to identify the clumps that are associated with JQ and remove other foreground clumps unassociated with the JQ galaxies. The Bagpipes configuration is identical to the one used for the photometric redshift estimation of ELG2, ELG3, and ELG4. We visually inspect the best-fit results and define a clump as a member of JQ if its best-fit photometric redshift lies within the range of 6.65 and 6.75. Once the member clumps of JQ are identified, we rerun the SED fitting with the redshift fixed to 6.71 to best model their properties assuming the spectroscopic redshift of ELG1 and ELG5. We note clumps C8-1 and C8-2 are heavily blended in the longer wavelength imaging data, so we combine their photometry for the SED fitting process. In Extended Data Table 1, we summarize the physical properties of the identified member clumps.
In Extended Data Figure 2, we compare the size – SFR relation of the clumps in the JQ system with those of the galaxies at similar redshifts observed with JWST NIRCam [22, 21, 23, 24]. Several studies [22, 21, 23] resolved the galaxies through the lensing technique and found that the star clusters and star-forming clumps identified within the individual galaxies have sizes pc and SFRs yr-1. Their sizes and SFRs are significantly smaller than those of the clumps in the JQ system. The JQ clumps have sizes and SFRs more similar to those unresolved galaxies at 5 – 14 [24], which typically have sizes pc and SFRs yr-1.
[O III]+H emission map of JQ
The [O III]+H emission map of JQ is created by subtracting the continuum contribution from the NIRCam F356W image. Previous works often use images at slightly longer or shorter wavelengths to estimate the continuum contribution, such as NIRCam F335M for our case. However, due to the NIRCam F335M image being much shallower than the F356W, subtracting F335M would significantly reduce the SNR, making the diffuse [O III]+H halo invisible.
In this work, we adopt an alternative method to estimate the continuum contribution. First, we remove the foreground galaxies using their best-fit flux in F356W and Sérsic models. We then estimate the continuum for clumps associated with JQ based on their best-fit spectra from the Bagpipes fitting. We mask the [O III] and H emission lines in the best-fit spectrum and convolve them with the transmission curve of the F356W filter to calculate their contribution to the F356W image. We then use their best-fit Sérsic models, the continuum flux density, and the F356W PSF to generate a continuum map (shown in Extended Data Figure 3b). We subtract this continuum map from the foreground-removed F356W image to generate the [O III]+H emission map, as shown in Extended Data Figure 3c.
Spatially-resolved photometry and properties
As aforementioned, we fix the half-light radii of clumps when fitting their morphologies in the F356W image. However, emission lines are frequently observed to be more spatially extended than the underlying stellar continuum [68], so this approach could miss diffuse [O III]+H emission from the clumps, underestimation of their flux densities. Moreover, the simple Sérsic profile is not always able to accurately describe the morphologies of the clumps due to their disturbed and irregular shapes. To address these limitations and provide a robust estimate of the total stellar mass and SFR of the JQ system, we perform spatially resolved SED fitting directly using multi-wavelength images. This method is insensitive to the assumption of morphologies across different wavelengths and has been extensively validated in previous studies [69]. For this analysis, we adopt the SED fitting package, CIGALE [70], which offers computational efficiency when fitting a large number of data points. This also allows us to assess the potential biases associated with the SED fitting methods.
Our method follows the procedure in ref [69]. In detail, we used the PSF-matched imaging with a pixel scale of 0.03 arcsec. To enhance the SNR of regions surrounding and connecting the galaxies in the JQ, we applied an adaptive binning algorithm222https://github.com/pierrethx/MVT-binning based on the Weighted Voronoi Tessellation method described in ref [71, 72]. We chose a minimum SNR of 8 in the F356W band, and a minimum Voronoi bin size of 9 pixels, which is approximately the size of the FWHM of the PSF of F356W (FWHM = 01).
We run CIGALE using the 9-band fluxes in each Voronoi bin. Here we described the modules and parameters adopted in CIGALE. We adopted a delayed SFH allowing the and stellar age to vary from 0.05 – 1 Gyr and 40 – 800 Myr, respectively. In addition to the main SFH, we allow for a more recent “burst” of star formation as an exponential SFH with a allowed to vary from 20 – 100 Myr, an age of 1 – 30 Myr, and a mass fraction of this burst () to vary from 0 – 0.9. We assumed a ref [73] IMF and the stellar population synthesis models presented by ref [74] with metallicity ranging from Solar () to sub-Solar values (0.4 and 0.2 ). We include nebular emission using templates of ref [75]. We allow the ionization parameter to vary between to , the gas metallicity () to vary between 0.002 and 0.02, and a fixed electron density of 100 cm-3. The dust attenuation follows the Calzetti extinction law [63] allowing the dust attenuation in emission lines from nebular regions to vary from 0 to 1.1, and a lower dust attenuation in stellar continuum with a fixed dust attenuation ratio () between emission lines and stellar continuum. In Extended Data Figure 4, we present the spatially resolved maps of mass and SFR.
The integrated properties of the JQ system are derived by summing the spatially resolved maps. We obtain a total stellar mass of and a total SFR averaged of yr-1. We estimate the uncertainties in the total stellar mass and SFR by propagating the measurement errors from each Voronoi bin. The total stellar mass is consistent with the sum of the stellar masses of individual clumps from Bagpipes, however, the total SFR is nearly double the sum of the SFRs of the clumps (see Extended Data Table 1). To investigate this discrepancy, we further use CIGALE to model the SEDs of 17 clumps and find that the estimated stellar masses and SFRs agree with those derived from Bagpipes within uncertainties. Therefore, the difference between the SFRs estimated from the spatially resolved map and the individual clumps is not due to the choice of the SED fitting method. These results imply that while the majority of stellar mass is formed within the clumps, a large fraction, , of the total SFR may arise from diffuse emission surrounding the clumps if attributing the diffuse emission to star formation.
Contextualizing JQ-like systems with cosmological simulations
We compare the presence and properties of the JQ system with galaxy-size clumps found in a simulated lightcone [26]. The simulated lightcone spans 2 deg2, containing simulated galaxies arranged in a way that mimics how we observe the real Universe. The dataset provides galaxy coordinates (RA and Dec) on the projected plane, and includes both cosmological redshifts and peculiar velocities of galaxies along the line of sight. This enables a direct comparison between the simulation and observation. The large volume of the lightcone enables more precise quantification of the number of JQ-like systems in the Universe, although this is significantly larger than the 120 arcmin2 covered by the JADES dataset. The lightcone is constructed by drawing halos from the SMDPL simulation from the MultiDark suite [76] at various redshifts based on their line-of-sight distances. Within these dark matter halos, galaxies are then simulated using the Santa Cruz semi-analytic model for galaxy formation [77, 78, 79, 80], which incorporates the evolution of a wide variety of physical processes, such as cosmological accretion, cooling, star formation, chemical enrichment, and stellar and AGN feedback. The performance of the model at high redshift is validated against one-point distribution functions of , , and SFR [79, 81] and galaxy clustering constraints [80]. We also note that dust attenuation is accounted for in the predicted photometry of simulated galaxies. Sky coordinates of the halos and galaxies therein are computed based on the physical coordinates of sources within the lightcone. We refer the readers to refs. [80, 26] for detailed descriptions for the simulation and the construction of the lightcone.
While the Santa Cruz SAM tracks the decay of satellite orbits due to dynamical friction and the tidal disruption of infalling satellites (see ref [82], for details), the model does not explicitly model star-forming clumps nor clumps that form during mergers. For comparison with the lightcone, we only consider the galaxy-size clumps in JQ that are distinctly separated to be the analogs to the galaxies in the lightcone. These include ELG1 (C1+C2+C3), ELG2 (C5+C6), ELG3 (C8+C9), ELG4 (C10), ELG5 (C11), C4, C7, C12+C13, C14, C16. We exclude C15 because it is relatively faint compared to the others. From the lightcone, we first create a parent galaxy sample at – 7.0 with F277W magnitude and then identify a subset of galaxy systems in the parent sample by requiring at least one companion with a velocity difference of km s-1 and a projected separation of arcsec. These criteria are chosen to be consistent with the observation of JQ and the depth of the JADES survey. The redshift range (6.5 – 7.0) corresponds to a comoving volume of cMpc3 for the lightcone simulation. In total, we select 29,696 unique systems (central and satellite galaxies) from the 2 deg2 lightcone, of which 3154 systems have at least two galaxy-size clumps (inferred by simulated effective radii and SFRs based on results presented in Ref [24]). This implies that we would expect galaxy systems with galaxy-size clumps at – 7.6 for the JADES survey area. In Extended Data Figure 5, we present the expected distribution of clump numbers in galaxy systems in the JADES survey. Using this criterion (systems with clumps), we identify 8 such systems in the 2 deg2 lightcone, corresponding to an expectation of just 0.26 systems within the JADES survey volume. This indicates the galaxy-size clump density in JQ is higher than galaxy systems in cosmological simulation.
The JQ system is serendipitously discovered in an ELG study based on the JADES F410M image. Given the relatively small survey volume of the JADES F410M image ( arcsec2), we estimate the observed comoving number density of the JQ system to be cMpc-3. The error is estimated based on the Poisson error of the small number statistic [83]. In comparison, we estimate the comoving number density of JQ-like systems in lightcone simulation to be cMpc-3.
Merging timescale of the JQ system
Constraining the merging timescale of galaxies is extremely challenging due to significant uncertainties, such as the degeneracy between line-of-sight distance and velocity and the unknown transverse velocity of the galaxies. In systems like JQ, the presence of multiple clumps further complicates precise calculations. To simplify the estimation, we only consider the merging of two main components in JQ. We define the primary component of JQ as the subsystem of ELG3 and ELG4, the most massive members of JQ. We define the secondary component as ELG1, the secondary massive galaxy, and the most distant in the projected plane. The projected distance between the primary and secondary galaxies is measured based on the projected distance between their most massive clumps, yielding a value of 13.3 kpc. The total stellar mass is determined by summing the stellar masses of the member clumps. We obtain the stellar mass of for the secondary and for the primary galaxy. To estimate the halo properties of the primary and secondary galaxies, we select the simulated galaxies at – 7.0 from the lightcone catalog [26] with stellar masses similar to those of our objects. We identify halos with mass ratios between the primary and secondary galaxies therein to be . The halo properties are determined by the median values of those simulated galaxies, with uncertainties defined by the 16th to 84th percentile from the distribution. The virial radius and velocity are kpc and km s-1 for the primary galaxy and kpc and km s-1 for the secondary galaxy.
We estimate the merging timescales in three ways. We first estimate the lower limit of the merging timescale, assuming the secondary galaxy follows a circular orbit about the primary galaxy, where dynamical friction dominates its energy and angular momentum losses. The merging timescale follows the formula (Equation 12.45 in ref [84]) based on Chandrasekhar’s dynamical friction formula [85],
where is the mass ratio between the primary and secondary galaxies, is the circular velocity, is the distance between the two galaxies, and is the virial radius of the primary galaxy. Adopting the virial velocity of the primary galaxy to be and the projected distance between the two galaxies to be , we obtain a merging timescale of Myr. It is important to note that is the projected distance and only represents a lower limit of the true spatial separation. If assuming a random orientation of primary and secondary galaxies, the expected deprojected distance would be kpc, and the merging timescale increases to Myr.
Second, equation 1 can underestimate the merging timescale because, during the merger, the tidal forces of the primary galaxy would gradually strip the matter from the secondary galaxy, decreasing its mass [86]. This effect would increase the merging timescale by a factor of 2 – 5 [87, 84]. To incorporate the effect of mass loss, we adopt the formula by Ref [88] (Equation 8.17),
where and are the circular velocity of the two galaxies. Adopting the virial velocities of the galaxies as the circular velocities and the projected distance between the two galaxies to be , we obtain a merging timescale of Myr. If considering the deprojected distance of kpc, the merging timescale increases to Myr.
Third, we adopt a the formula of ref [87] (Equation 8), which fits results from an -body numerical simulations, to estimate the merging timescale:
where is a constant, approximately 0.43, and is the circularity parameter of the secondary’s orbit. Similar to the previous two models, we assume a circular orbit () and adopt the virial velocity as . We obtain a merging timescale of (320) Myr for a distance of 13.3 (16.3) kpc.
Therefore, we conclude the merging timescale of JQ is 100 – 350 Myr, which spans the range from all the estimates above.
No evidence of outflows in ELG1
In the bottom right panel of Extended Data Figure 6, we present the asymmetric [O III] emission line of ELG1 from the NIRSpec Medium Grating observation. This asymmetric profile does not indicate the presence of galactic outflows, instead, it is due to the two clumps (C1 and C2) being simultaneously targeted within the same NIRSpec shutter, as shown in Extended Data Figure 6a. The alignment between C1 and C2 is nearly parallel to the dispersion direction, with a separation of arcsec. This implies a velocity offset of km s-1 between C1 and C2 in the dispersion direction of the NIRSpec spectrum. To model their emission line profile, we fit their [O III] line with two Gaussian components. The centers of two components are linked based on a velocity difference of km s-1. We find that the observed [O III] profile can be well reproduced by only these two components with widths of km s-1 and km s-1, significantly narrower than the typical outflow width ( km s-1) observed in high-redshift galaxies [89] and their local analogs [90]. The residual (gray dotted line) is consistent with zero, confirming that the line asymmetry is solely due to the superposition of the two clumps. Therefore, this suggests that ELG1 lacks strong outflows.
Can outflows explain the [O III]+H bridge between ELG1 and ELG2?
Although the JWST NIRSpec spectrum of ELG1 does not reveal evidence of outflows, this does not conclusively rule out the presence of outflows in other clumps within ELG1 or in ELG2, which were not targeted by the NIRSpec observations. In this section, we further explore whether the outflows in ELG1 and ELG2 and their adjacent clumps can produce the [O III]+H bridge between them based on the outflow simulations. We estimate the [O III]+H luminosity of [O III]+H bridge to be the total flux enclosed by a polygon region defined in Extended Data Figure 7 with the contribution of clumps C12 and C13 removed based on Extended Data Table 1. We derive the total observed [O III]+H luminosity of the bridge to be erg s-1.
We first consider a single-phase galactic wind model, where the direct radiative cooling of hot winds from K to K generates the emission lines [91, 90]. Under this model, we calculate the extended [O III]+H emission produced by ELG1 and ELG2. We consider supernova explosions the dominant sources of mass and energy in the hot winds. Therefore, the energy and mass input rates are proportional to the SFR:
where and are the thermalization-efficiency and mass-loading factors of hot winds, respectively. Measuring these factors is extremely challenging because the hot winds can only be observed in the X-ray. Here, we consider 0.1 and 1 for and consider a wide range of 0.2 – 1.6 with a step of 0.2 for . These values are consistent with those derived from numerical simulations [92, 93, 94]. The chosen values are also consistent with the X-ray observation of hot winds in the nearby starburst galaxy M82 [95]. For the input SFRs, we adopt those measured from the SED fitting for individual clumps. We assume a spherical geometry where the galactic winds launch at the half-light radius of galaxies (0.4 – 0.6 kpc) and expand isotropically. The radiative cooling of the hot winds follows the cooling function at in ref [96]. The studies of local analogs of high-redshift galaxies revealed that their galactic outflows are usually more metal-enriched [97] than the ISM. Thus, in this work, we adopt the metallicity of hot winds to be the solar metallicity. We utilize the escape velocity to quantify the impact of gravity on the outflow. The escape velocity relies on the radial distribution of dark matter and is typically larger than the circular velocity by a factor of [88]. Therefore, we adopt a wide range of 200 – 500 km s-1 with a step of 100 km s-1 to account for those various uncertainties. We integrate the luminosities of [O III] doublets and H emission lines from the launch radius to 8 kpc, where the latter corresponds to the distance between ELG1 and ELG2. We find the maximal [O III]+H luminosity (when ) is only erg s-1, 2 – 3 orders of magnitude lower than the luminosity of [O III]+H bridge between ELG1 and ELG2.
Recent studies have revealed the multiphase nature of galactic outflows [98, 99]. The temperature and shear velocity differences between the hot and cold ( K) phases drive the mixing of the two phases and generate the turbulent radiative mixing layers (TRMLs), where the temperature drops into the intermediate regime ( K) [100, 101]. This intermediate temperature regime corresponds to the peak of the radiative cooling function, facilitating the production of [O III] and H emission lines. Nevertheless, Ref [90] found that the direct radiative cooling flux of hot winds becomes comparable to () that of TRMLs when . Additionally, due to the complex interplay between radiative cooling, conduction, and viscous heating in the TRMLs, the cooling flux of TRMLs peaks at K, where the fraction of radiative cooling contributes to [O III] becomes relatively small compared to that of hot winds (ratio . As a result of these two factors, the multi-phase galactic wind model with TRMLs [101] can only enhance the total [O III]+H luminosity by dex. This remains insufficient to explain the luminosity of the [O III]+H bridge between ELG1 and ELG2.
Consequently, galactic outflows are unlikely to be the dominant drivers of [O III]+H bridge. Therefore we favor the scenario where the [O III]+H bridge originates from shock-heated, tidally stripped gas.
Correspondence and request for materials should be addressed to W. Hu.
Acknowledgments This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). This work is based on observations made with the NASA/ESA/CSA JWST. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. This work was supported in part by generous funding provided by Marsha and Ralph Schilling through Texas A&M University.
Author contribution W.H. and C.P. designed the layout of this paper. W.H. reduced the data, performed scientific analysis, and wrote the manuscript. C.P. helped with the manuscript writing and scientific analysis. L.S. helped with the spatially resolved analysis. C.P. and L.Y.A.Y. helped with the lightcone simulation and merging timescale estimation. Z.P. performed the outflow analysis based on analytical galactic wind models. All authors discussed the results and commented on the manuscript.
Conflict of interest/Competing interests The authors declare no competing interests.
Data availability The unprocessed JWST data are available through the Mikulski Archive for Space Telescopes (https://mast.stsci.edu/search/ui/#/jwst). The reduced JWST images and spectra in this work are publicly available through the STSci High-Level Science Products (https://archive.stsci.edu/hlsp/jades). The mock catalogs from the lightcone simulation are publicly available through the Semi-analytic forecasts for the Universe homepage (https://www.simonsfoundation.org/semi-analytic-forecasts).
Code availability Codes used in this study are publicly available: Astropy (https://www.astropy.org), Bagpipes (https://github.com/ACCarnall/bagpipes), CIGALE (https://cigale.lam.fr), Pysersic (https://github.com/pysersic/pysersic), sep (https://github.com/sep-developers/sep), and SExtractor (https://www.astromatic.net/software/sextractor).